OpenStudy (anonymous):

can someone plz help me with this question....marcus wants to buy some candy. His favvourites are fruit blasters(x) and sour suckers(y). Fruit blasters cost 0.05 each and sour suckers 0.10. he buys a total of 20 candies for 1.20. determine the number of each candy he bought.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

x+y=20 0.05x + 0.10y = 1.20 just solve it

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

yya but help me with it

OpenStudy (anonymous):

plz and thnxxx

OpenStudy (anonymous):

ok x = 16 y = 4

OpenStudy (anonymous):

can you plz explain i have a test on it

OpenStudy (anonymous):

? u still there?

OpenStudy (anonymous):

see you can use the first equation and find x .... would be x = 20 - y then you substitute it in the second equation. or you can multiply the first equation by -0.05 then sum both, which you'll only have the y variable, you can find it, then you get that value back to the 1 equation

OpenStudy (anonymous):

you can do that but you can also use matrix and rref it, which I don't think you know how to do it right?

OpenStudy (anonymous):

oh thnx that really helps alot!

OpenStudy (anonymous):

i am grd 9 so not really

OpenStudy (anonymous):

wat about u?/

OpenStudy (anonymous):

I don't think you gonna learn it if you only takes regular classes, I don't know how the school works on USA.. I'm on engineering college on Brazil

OpenStudy (anonymous):

iam from canada

OpenStudy (anonymous):

oh wow u r a pro!!!

OpenStudy (anonymous):

I can handle things from school

OpenStudy (anonymous):

Awesome!

OpenStudy (anonymous):

send me a message if you need something if I know what it means I'll help

OpenStudy (anonymous):

honestly you really helped so thnx alot but how do i find u agian online.